OpenStudy (anonymous):

A building measures 50 feet in length and w feet in width. Which of the following is an expression for the perimeter of the building? A-50/w b-50w c-50+w d-100+2w

OpenStudy (anonymous):

d-100+2w

OpenStudy (turingtest):

P=2L+2w=2(50)+w=100+w

OpenStudy (anonymous):

@Turing it shud be 2w

OpenStudy (anonymous):

Perimeter = 50 + w + 50 + w = 100 + 2w

OpenStudy (turingtest):

True; typo.

OpenStudy (anonymous):

no problemo, It happenso.....☺